\begin{thebibliography}{99}
\bibitem{nbs}M. Abramowitz and I. A. Stegun,
{Handbook of Mathematical Functions, Washington, National Bureau of Standards
(1964).}
%\bibitem{apostol}T. M. Apostol,
%{Introduction to Analytic Number Theory, Springer Verlag, New York (1976).}
%\bibitem{balazard}M. Balazard, E. Saias, and M. Yor,
%{Notes sur la fonction $\zeta$ de Riemann, 2, Advances in Mathematics 
%{\bf 143}, 284-287 (1999).}
\bibitem{bala}U. Balakrishnan,
{On the Laurent expansion of $\zeta(s,a)$ at $s=1$, J. Indian Math. Soc.
{\bf 46}, 181-187 (1982).}
\bibitem{berndt}B. C. Berndt,
{On the Hurwitz zeta function, Rocky Mtn. J. Math. {\bf 2}, 151-157 (1972).}
%\bibitem{berry}M. V. Berry,
%{Quantum chaology, Proc. Roy. Soc. London A {\bf 413}, 183-198 (1987); M. V.
%Berry and J. P. Keating, A new asymptotic representation for $\zeta(1/2+it)$
%and quantum spectral determinants, Proc. Roy. Soc. London A {\bf 437}, 
%151-173 (1992).}
%\bibitem{askey}R. Askey and G. Gasper,
%{J. Anal. Math. {\bf 31}, 48-68 (1977).}
%\bibitem{backlund}R. J. Backlund,
%{\"{U}ber die Nullstellen der Riemannschen Zetafunktion, Acta Math. {\bf 41},
%345-375 (1918).}
%\bibitem{pfqref}W. N. Bailey,
%{Generalized Hypergeometric Series, Cambridge University Press (1935);
%G. E. Andrews, R. Askey, and R. Roy, Special Functions, Cambridge University
%Press (1999).}
\bibitem{biane}P. Biane, J. Pitman, and M. Yor,
{Probability laws related to the Jacobi theta and Riemann zeta functions and
Brownian excursions, Bull. Amer. Math. Soc. {\bf 38}, 435-465 (2001).} %  On 
%p. 441, the approximate numerical value of $k_4$ should read $k_4 \simeq
%-0.000441763$.  On p. 440, $E(Y^{1-s})=2\xi(1-s)=2\xi(s)$.}
\bibitem{bloch}S. Bloch,
{Zeta values and differential operators on the circle, J. Algebra {\bf 182},
476-500 (1996).}
\bibitem{briggs}W. E. Briggs, 
{Some constants associated with the Riemann zeta-function, 
Mich. Math. J. {\bf 3}, 117-121 (1955).}
%; J. Lepowsky, Vertex operator algebras and the 
%zeta function, arXiv:math.QA9909178 (1999); D. Kreimer, The residues of 
%quantum field theory--numbers we should know, arXiv:hep-th/04040909 (2004).}
%\bibitem{bohigas}O. Bohigas and M.-J. Giannoni,
%{in:  Lecture Notes in Physics, Vol. 209, Springer (1984), p. 1; O. Bohigas,
%M.-J. Giannoni, C. Schmit, Phys. Rev. Lett. {\bf 52}, 1 (1984); H. L. 
%Montgomery, Proc. Symp. Pure Math. {\bf 24}, 181 (1973).}
%\bibitem{bombieri}E. Bombieri and J. C. Lagarias,
%{Complements to Li's criterion for the Riemann hypothesis,
%J. Number Theory {\bf 77}, 274-287 (1999).}
%\bibitem{bump}D. Bump,
%{Automorphic Forms and Representations, Cambridge University Press (1997).}
%\bibitem{clark}W. E. Clark and M. E. H. Ismail,
%{Inequalities involving gamma and psi functions,
%http://www.math.usf.edu/\~ clark, to appear in Analysis and Applications.}
\bibitem{coffey}M. W. Coffey,
{Relations and positivity results for derivatives of the Riemann $\xi$
function, J. Comput. Appl. Math., {\bf 166}, 525-534 (2004).}
%\bibitem{cofjcam1}M. W. Coffey,
%{On some log-cosine integals related to $\zeta(2)$, $\zeta(3)$, and 
%$\zeta(6)$, to appear in J. Comput. Appl. Math.}
\bibitem{coffey03}M. W. Coffey,
{Toward verification of the Riemann hypothesis:  Application of the
Li criterion, to appear in Math. Phys., Analysis and Geometry (2005).}
\bibitem{coffeygam1}M. W. Coffey,
{New results on power series expansions of the Riemann xi function and 
the Li/Keiper constants, preprint, 2004.}
\bibitem{dilcher}K. Dilcher,
{Generalized Euler constants for arithmetical progressions, Math. Comp.
{\bf 59}, 259-282 (1992).}
%\bibitem{coffeykremin}M. W. Coffey,
%{New results on the Stieltjes constants:  Proof of a Kreminski conjecture, 
%preprint, 2004.}
%%\bibitem{coffey04}M. W. Coffey,
%%{The Li constants are asymptotically positive, and more, preprint, 2004.}
%\bibitem{coffeyhecke}M. W. Coffey,
%{An explicit formula and estimations for Hecke L functions, preprint, 2004.}
%\bibitem{cofunpub}M. W. Coffey,
%{unpublished (2004).}
%\bibitem{cofpla}M. W. Coffey,
%{A set of identities for a theta function at unit argument, Phys. Lett. A,
%{\bf 300}, 367-369 (2002).}
%\bibitem{davenport}H. Davenport,
%{Multiplicative Number Theory, Springer Verlag (2000).}
%\bibitem{dunster}T. M. Dunster,
%{Uniform asymptotic expansions for Whittaker's confluent hypergeometric
%functions, SIAM J. Math. Anal. {\bf 20}, 744-760 (1989).}
\bibitem{edwards}H. M. Edwards,
{Riemann's Zeta Function, Academic Press, New York (1974).}
\bibitem{el}E. Elizalde, 
{Ten physical applications of spectral zeta functions, Springer (1995).}
%\bibitem{frenzen}C. L. Frenzen and R. Wong,
%{Uniform asymptotic expansions of Laguerre polynomials, SIAM J. Math. Anal.
%{\bf 19}, 1232-1248 (1988).}
%\bibitem{goldfeld}D. Goldfeld,
%{Explicit formulae as trace formulae, in:  Number Theory, Trace formulas, and 
%Discrete Groups, pp. 281-288, Academic Press (1989).}
\bibitem{grad}I. S. Gradshteyn and I. M. Ryzhik,
{Table of Integrals, Series, and Products, Academic Press, New York (1980).}
%\bibitem{graham}R. L. Graham, D. E. Knuth, and O. Patashnik,
%{Concrete Mathematics, 2nd ed., Addison Wesley (1994).}  %look up this book 
%\bibitem{guinand}A. P. Guinand,
%{A summation formulae in the theory of prime numbers, Proc. London Math. Soc.
%(2) {\bf 50}, 107-119 (1948).}
\bibitem{hansen}E. R. Hansen and M. L. Patrick,
{Some relations and values for the generalized Riemann zeta function,
Math. Comp. {\bf 16}, 265-274 (1962).}
\bibitem{hasse}H. Hasse,
{Ein summierungsverfahren f\"{u}r die Riemannsche Zeta-Reihe, Math. Z. 
{\bf 32}, 458-464 (1930).}
%\bibitem{havil}J. Havil,
%{Gamma:  Exploring Euler's Constant, Princeton University Press (2003).}
%\bibitem{hawking}S. W. Hawking,
%{Zeta function regularization of path integrals in curved spacetime, Commun. 
%Math. Phys. {\bf 55}, 133-148 (1977).}
%\bibitem{hua}L. K. Hua,
%{Introduction to Number Theory, Springer Verlag (1982).}
%\bibitem{scref}M. I. Israilov,
%{On the Laurent Decomposition of Riemann's zeta function,
%Dokl. Akad. Nauk SSSR (Russian) {\bf 12}, 9 (1979); M. I. Israilov,
%Trudy Mat. Inst. Steklova {\bf 158}, 98-104 (1981); G. H. Hardy, 
%Note on Dr. Vacca's series for $\gamma$, Quart. J. Pure
%Appl. Math. {\bf 43}, 215-216 (1912); J. C. Kluyver, On certain series of
%Mr. Hardy, Quart. J. Pure Appl. Math. {\bf 50}, 185-192 (1927); W. E. 
%Briggs, Some constants associated with the Riemann zeta-function, 
%Mich. Math. J. {\bf 3}, 117-121 (1955); D. Mitrovi\'{c}, The signs of some
%constants associated with the Riemann zeta function, Mich. Math. J. {\bf 9}, 
%395-397 (1962).} 
\bibitem{ivic}A. Ivi\'{c}, 
{The Riemann Zeta-Function, Wiley (1985).}
%\bibitem{ivic93}A. Ivi\'{c},
%{The Laurent coefficients of certain Dirichlet series,
%Publs. Inst. Math. {\bf 53}, 23-36 (1993).  In Eq. (1.7), $(\log x)^x$ should
%be replaced with $(\log x)^k$ and on p. 25 "comparing with (1.15)" should be 
%replaced with "comparing with (1.5)".}
\bibitem{kanemitsu}S. Kanemitsu,
{On evaluation of certain limits in closed form, Th\'{e}orie des nombres,
J.-M. de Koninck and C. Levesque, eds., Walter de Gruyter, pp. 459-474 (1989).}
\bibitem{karatsuba}A. A. Karatsuba and S. M. Voronin,
{The Riemann Zeta-Function, Walter de Gruyter, New York (1992).}
\bibitem{keiper}J. B. Keiper,
{Power series expansions of Riemann's $\xi$ function, Math. Comp. {\bf 58}, 
765-773 (1992).}
\bibitem{kluyver}J. C. Kluyver, 
{On certain series of Mr. Hardy, Quart. J. Pure Appl. Math. {\bf 50}, 
185-192 (1927).}
%\bibitem{knopp}K. Knopp,
%{Theory and Application of Infinite Series, Hafner Publishing Company, New
%York (1951).}
%\bibitem{katz}N. M. Katz and P. Sarnak,
%{Zeros of zeta functions and symmetry, Bull. Amer. Math. Soc., {\bf 36},
%1-26 (1999).}
%\bibitem{koepf}W. Koepf and D. Schmersau,
%{Bounded nonvanishing functions and Bateman functions, Complex Variables, 
%{\bf 25}, 237-259 (1994).}
\bibitem{kreminski}R. Kreminski,
{Newton-Cotes integration for approximating Stieltjes (generalized Euler)
constants, Math. Comp. {\bf 72}, 1379-1397 (2003).}  
%\bibitem{kreminski}R. Kreminski,
%{private communication (2004).}
%\bibitem{lachaud}G. Lachaud,
%{Spectral analysis and the Riemann hypothesis, J. Comp. Appl. Math., {\bf 160},
%175-190 (2003).  On p. 176 of this paper, $\xi(s)$ is missing the factors
%$s(s-1)/2$ and on p. 178 the definition of ${\cal L}$ is misstated.}
%\bibitem{leboeuf}P. Leboeuf and A. G. Monastra,
%{Quantum thermodynamic fluctuations of a chaotic Fermi-gas model, 
%Nucl. Phys. A, {\bf 724}, 69-84 (2003).}
%\bibitem{knuth}D. E. Knuth,
%{The Art of Computer Programming, Vol. 3, Addison Wesley (1973).}  %look up 
%\bibitem{lehmer}D. H. Lehmer,
%{The sum of like powers of the zeros of the Riemann zeta function, Math.
%Comput. {\bf 50}, 265-273 (1988).}
%\bibitem{lehmer}D. H. Lehmer,
%{The sum of like powers of the zeros of the Riemann zeta function, Math. Comp.
%{\bf 50}, 265-273 (1988).}
\bibitem{li}X.-J. Li,
{The positivity of a sequence of numbers and the Riemann hypothesis,
J. Number Th. {\bf 65}, 325-333 (1997).}
%\bibitem{li2}X.-J. Li,
%{Explicit formulas for Dirichlet and Hecke L-functions, Ill. J. Math., to 
%appear (2004).}
%\bibitem{li2001}X.-J. Li,
%{A formula for the Dedekind $\xi$-function of an imaginary quadratic field, 
%J. Math. Anal. Appl. {\bf 260}, 404-420 (2001).}
%\bibitem{luke}Y. L. Luke,
%{The Special Functions and Their Approximations, Academic Press (1969).}
%\bibitem{krzysztof}K. Ma\'{s}lanka,
%{http://nac.oa.uj.edu.pl/\~maslanka and Some comments concerning Li's criterion
%for Riemann Hypothesis, preprint (2003)}
\bibitem{mitrovic}D. Mitrovi\'{c}, 
{The signs of some constants associated with the Riemann zeta function, 
Mich. Math. J. {\bf 9}, 395-397 (1962).} 
%\bibitem{hlmont}H. L. Montgomery and A. M. Odlyzko,
%{Large deviations of sums of independent random variables, Acta Arith.
%{\bf XLIX}, 427-434 (1988).}
%\bibitem{mati}Yu. V. Matiyasevich,
%{A relationship between certain sums over trivial and nontrivial zeros of the
%Riemann zeta function, Mat. Zametki {\bf 45}, 65-70 (1989).}
%\bibitem{matsuoka}Y. Matsuoka,
%{A note on the relation between generalized Euler constants and the zeros of the
%Riemann zeta function, J. Fac. Educ. Shinshu Univ. {\bf 53}, 81-82 (1985);
%A sequence associated with the zeros of the Riemann zeta function, Tsukuba J.
%Math. {\bf 10}, 249-254 (1986).}
%\bibitem{matsuoka2}Y. Matsuoka,
%{Generalized Euler constants associated with the Riemann zeta function,
%in:  Number Theory and Combinatorics, ed. by J. Akiyama et al., World 
%Scientific, pp. 279-295 (1985).  Concerning just Section 6 here, $t$ should 
%read $T$ on p. 293 (line 4), $\Gamma$ should read $T$ on the same line, and 
%$+$ should read $=$ on the next line.  On the next page, $\gamma_0^3$ should 
%be replaced with $\gamma_0^4$ in Example 1.}
%\bibitem{mestre}J.-F. Mestre,
%{Formules explicites et minorations de conducteurs de vari\'{e}t\'{e}s 
%alg\'{e}briques, Compositio Math. {\bf 58}, 209-232 (1986).}
%\bibitem{cofpla02}M. W. Coffey,
%{Phys. Lett. A {\bf 300}, 367 (2002).}
%\bibitem{neukirch}J. Neukirch,
%{Algebraic Number Theory, Springer Verlag (1999).}
%\bibitem{odlyzko}A. M. Odlyzko,
%{http://www.dtc.umn.edu/\~odlyzko/} %put something else in for this reference?
%\bibitem{amo1989}A. M. Odlyzko,
%{Bounds for discriminants and related estimates for class numbers, regulators 
%and zeros of zeta functions:  A survey of recent results, Sem. Th. No., 
%Bordeaux, {\bf g}, 1-15 (1989).}
%\bibitem{omar}S. Omar,
%{Localization of the first zero of the Dedekind zeta function, Math. Comp.
%{\bf 70}, 1607-1616 (2001).}
%\bibitem{rao}K. S. Rao, G. V. Berghe, and C. Krattenthaler,
%{An entry of Ramanujan on hypergeometric series in his notebooks, preprint
%arXiv:Math.CA/0304317 v1 (2003).}
%\bibitem{other}A. B. Olde Daalhuis and N. M. Temme,
%{Uniform Airy-Type Expansions of Integrals, SIAM J. Math. Anal. {\bf 25}, 
%304-321 (1994).}
%\bibitem{olver}F. W. J. Olver,
%{Asymptotics and Special Functions, Academic Press (1974).}
%\bibitem{planat}M. Planat,
%{1/f noise, the measurement of time and number theory, Fluc. Noise Lett.
%{\bf 1}, R65-R79 (2001).}
\bibitem{riemann}B. Riemann,
{\"{U}ber die Anzahl der Primzahlen unter einer gegebenen Gr\"{o}sse, 
Monats. Preuss. Akad. Wiss., 671 (1859-1860).}
%\bibitem{rooney}P. G. Rooney,
%{Further inequalities for generalized Laguerre polynomials, C. R. Math. Rep.
%Acad. Sci. Canada, {\bf 7}, 273-275 (1985).}  %can get without ordering in mail?
\bibitem{sondow}J. Sondow,
{Analytic continuation of Riemann's zeta function and values at negative integers 
via Euler's transformation, Proc. Amer. Math. Soc. {\bf 120}, 421-424 (1994).}
\bibitem{sri}H. M. Srivastava and J. Choi,
{Series associated with the zeta and related functions, Kluwer (2001).}
\bibitem{stieltjes}T. J. Stieltjes,
{Correspondance d'Hermite et de Stieltjes, Volumes 1 and 2, Gauthier-Villars,
Paris (1905).}
%\bibitem{szego}G. Szeg\"{o},
%{Orthogonal Polynomials, Vol. 23 of AMS Colloquium Publications, AMS (1975).}
%look up this reference:
%\bibitem{temme}N. Temme,
%{Asymptotic estimates for Laguerre polynomials, J. Appl. Math. Phys. {\bf 41},
%114-126 (1990).}  %also to look up:
\bibitem{titch}E. C. Titchmarsh,
{The Theory of the Riemann Zeta-Function, 2nd ed., Oxford University
Press, Oxford (1986).}
%\bibitem{voros}A. Voros,
%{A sharpening of Li's criterion for the Riemann hypothesis, arXiv:math.NT/0404213
%(2004); Zeta functions for the Riemann zeros, Ann. Inst. Fourier, Grenoble {\bf 
%53}, 665-699 (2003).}
%\bibitem{watson}E. T. Whittaker and G. N. Watson,
%{A course of modern analysis, 4th ed., Cambridge Univ. Press, Cambridge (1962).}
%\bibitem{weil}A. Weil,
%{Basic Number Theory, Springer Verlag (1967).}
%\bibitem{weil}A. Weil,
%{Oeuvres Scientifiques--Collected Papers, Vol. II, pp. 48-61, Springer
%(1980).}
\bibitem{wilton}J. R. Wilton, 
{A note on the coefficients in the expansion of $\zeta(s,x)$ in powers of 
$s-1$, Quart. J. Pure Appl. Math. {\bf 50}, 329-332 (1927).}
%\bibitem{wu}H. Wu and D. W. L. Sprung,
%{Riemann zeros and a fractal potential, Phys. Rev. E, {\bf 48}, 2595-2598 
%(1993).}
\bibitem{zhang}N.-Y. Zhang and K. S. Williams,
{Some results on the generalized Stieltjes constants, Analysis {\bf 14}, 
147-162 (1994).}
%In addition to a typographical error present in the proof
%of Theorem 5, Remark 5, and Eq. (1.11) of this reference, where $2^{-n}$
%should read $2^{-n-1}$, the first term on the right side of Eq. (6.2) should 
%read $(\log^n a)/a$, and in both Eqs. (6.5) and (6.11), $P_n$ should appear in 
%place of $P_1$.  On p. 148, in Eq. (1.4), $f_n(x)$ should be replaced by
%$f_n'(x)$.  On p. 157, (7.2) should be replaced with 
%(7.1) in the second line of text from the bottom, and on p. 158 (7.2) should 
%be replaced with (1.9) in the third line of text.}  %from the top.}
\end{thebibliography}

\documentstyle[12pt]{article}
\title{New results on the Stieltjes constants:  Asymptotic and
exact evaluation} 
\author{Mark W. Coffey\\
Department of Physics\\
Colorado School of Mines\\
Golden, CO  80401\\
(Received $\mbox{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~2005}$)}
\date{April 9, 2005}
\begin{document}
\maketitle
\baselineskip=25 pt
\begin{abstract}

The Stieltjes constants $\gamma_k(a)$ are the expansion coefficients in
the Laurent series for the Hurwitz zeta function about $s=1$.
We present new asymptotic, summatory, and other exact expressions for 
these and related constants.  

\end{abstract}
 
\vspace{.25cm}
\baselineskip=15pt
\centerline{\bf Key words and phrases}
\medskip 

\noindent

Stieltjes constants, Riemann zeta function, Hurwitz zeta function, Laurent 
expansion, integrals of periodic Bernoulli polynomials, functional equation,
Kreminski conjecture

 
\baselineskip=25pt
\pagebreak
\medskip
\centerline{\bf Introduction}
\medskip

The Stieltjes (or generalized Euler) constants 
\cite{bala,briggs,dilcher,ivic,kluyver,kreminski,mitrovic,stieltjes,wilton,zhang}
$\gamma_k(a)$ appear as 
expansion coefficients in the Laurent series about $s=1$ of the Hurwitz 
zeta function $\zeta(s,a)$, a generalization of the Riemann zeta 
function $\zeta(s)$ \cite{berndt,edwards,ivic,karatsuba,riemann,titch}.
We present both new asymptotic and 
exact expressions for these and other fundamental mathematical constants.  

In the following, we let $s(n,m)$ denote the Stirling numbers of the first
kind and $C_k(a) \equiv \gamma_k(a) -(\ln^k a)/a$.  By convention, 
$\gamma_k$ represents $\gamma_k(1)$.  With $B_n(x)$ the
Bernoulli polynomials, their periodic extension is denoted 
$P_n(x) \equiv B_n(x-[x])$.  As is customary, $B_j$ denotes the Bernoulli
numbers $B_j(0)=(-1)^jB_j(1)$.

{\bf Proposition 1}.  As $n \to \infty$, we have $C_n(a+1/2) \to -C_n(a)$.

{\bf Proposition 2}.  For integers $m>0$ and $n \geq 1$ we have 
$$\gamma_n=(-1)^{n-1}n!m^{1-n}\sum_{k=0}^{n+1}{{s(n+1,n+1-k)} \over {k!}}
\int_1^\infty P_n(mx) {{\ln^k x} \over x^{n+1}} dx
-\sum_{r=1}^{m-1} C_n(r/m).$$

{\bf Proposition 3}.  For integers $q \geq 2$ we have
$$\sum_{r=1}^{q-1} \gamma_k(r/q)=-\gamma_k + q(-1)^k {{\ln^{k+1} q} \over
{(k+1)}} + q\sum_{j=0}^k {k \choose j} (-1)^j (\ln^j q) \gamma_{k-j}.$$

{\bf Proposition 4}.  For $\sigma = $ Re $s<1$ we have
$$\sum_{k=0}^\infty {{(-1)^k} \over k!} (s-1)^k \int_0^1 \gamma_k(a) da
={1 \over {1-s}}, ~~~~~~ \sigma <1.$$

{\bf Proposition 5}.  For large but fixed $k$, $0 < a \leq 1$, we have
$$\gamma_k(a) \sim m \sin[2\pi(a+\phi)],$$
for real $m$ and $0 \leq \phi <1$.

{\bf Proposition 6}.  For integers $n \geq 1$ and $k \geq 1$ we have  
$$\sum_{j=0}^{n-1} C_k\left(a \pm {j \over n}\right )=(-1)^{k-1}k!n^{1-k}
\sum_{\ell=0}^{k+1} {{s(k+1,k+1-\ell)} \over {\ell!}}\int_1^\infty P_k[n(x-a)] 
{{\ln^\ell x} \over x^{k+1}} dx.$$

{\bf Proposition 7}.  For integers $j \geq 1$ we have the set of coupled
differential equations
$${{(-1)^j} \over {j!}}{{d\gamma_j(a)} \over {da}}=-\sum_{k=j-1}^\infty
{{(-1)^k} \over {k!}}{{k+1} \choose j}\gamma_k(a), ~~~~~~j \geq 1,$$
and
$$-\psi'(a)=-1-\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\gamma_k(a),$$
where $\psi$ is the digamma function.

{\bf Proposition 8}.  We have (a)
$$\gamma_0=\gamma={1 \over 2}\ln 2-{1 \over {\ln 2}}
\sum_{n=0}^\infty {1 \over 2^{n+1}}\sum_{k=1}^n (-1)^k{n \choose k} {{\ln(k+1)} 
\over {(k+1)}},$$ 
for the Euler constant, (b)
$$-\gamma_1= {{\ln^2 2} \over {12}}-{1 \over 2}
\sum_{n=0}^\infty {1 \over 2^{n+1}}\sum_{k=1}^n (-1)^k{n \choose k} {{\ln(k+1)} 
\over {(k+1)}}+{1 \over {2\ln 2}}\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=1}^n (-1)^k{n \choose k} {{\ln^2(k+1)} \over {(k+1)}},$$
(c)
$$\eta_1={{\ln^2 2} \over {12}}+{1 \over 2}{1 \over {\ln^2 2}}
\left[\sum_{n=0}^\infty {1 \over 2^{n+1}}\sum_{k=1}^n (-1)^k{n \choose k} 
{{\ln(k+1)} \over {(k+1)}}\right]^2-{1 \over {\ln 2}}\sum_{n=0}^\infty 
{1 \over 2^{n+1}}\sum_{k=1}^n (-1)^k{n \choose k} {{\ln^2(k+1)} \over 
{(k+1)}},$$
(d)
$$\ln \pi = \ln 2 - 2\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=1}^n (-1)^k {n \choose k} \ln(k+1),$$
(e)
$$(-1)^{n-1}{{2^{2n-1}\pi^{2n}} \over {(2n)!}}B_{2n}=
{1 \over {1-2^{1-2n}}}\sum_{\ell=0}^\infty {1 \over 2^{\ell+1}}
\sum_{k=0}^\ell (-1)^k {\ell \choose k} {1 \over {(k+1)^{2n}}}$$
$$={1 \over {(2n-1)}}\sum_{\ell=0}^\infty {1 \over {\ell+1}}
\sum_{k=0}^\ell (-1)^k {\ell \choose k} {1 \over {(k+1)^{2n-1}}},
~~~~~~ n \neq 0,$$
and (f) the summation identity
$${1 \over {(2^{2n+1}-1)}}\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=1}^n (-1)^k {n \choose k} \ln(k+1) (k+1)^{2n}$$
$$=(-1)^n {{(2n)!} \over {2(2\pi)^{2n}}}{1 \over {(1-2^{-2n})}}
\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=0}^n (-1)^k {n \choose k} {1 \over {(k+1)^{2n+1}}}.$$
In part (c), the constants $\eta_j$ are 
defined by the expansion \cite{coffey,coffey03,coffeygam1}
$\ln \zeta(s)=-\ln(s-1)-\sum_{p=1}^\infty {\eta_{p-1} \over p}(s-1)^p$, 
$s \neq 1$.

These theorems are representative, but not comprehensive, of our results.
We do not reproduce relevant background material on properties and
functional equations of the special functions involved, but simply mention 
the useful sources \cite{nbs,grad,hansen,sri,zhang}.

\centerline{\bf Proofs of the Propositions}
\medskip

Here we summarize or otherwise simply indicate the method of proof of the
Propositions above.

Proposition 1.  A key observation is that
$$P_{2n}(x-a)=(-1)^{n-1}{{2(2n)!} \over {(2\pi)^{2n}}}\left[\cos 2\pi (x-a)
+ O(2^{-2n}) \right ], \eqno(1)$$
and similarly for $P_{2n+1}(x-a)$, based upon the Fourier expansions of
these polynomials \cite{nbs}.  So for large $n$, $P_{2n}(x-a) = -P_{2n}
(x-a-1/2)+O(2^{-2n})$ and we may appeal to the representation
\cite{zhang}
$$C_n(a) = (-1)^{n-1} n!\sum_{k=0}^{n+1} {{s(n+1,n+1-k)} \over {k!}}
\int_1^\infty P_n(x-a) {{\ln^k x} \over x^{n+1}} dx, ~~n \geq 1. \eqno(2)$$
Due to the boundedness of $P_n$, $|P_n(x)| \leq [3+(-1)^n]/(2\pi)^n$ for
$n \geq 1$ \cite{berndt}, there is uniform convergence of the integral in
this equation.  Therefore, we may use expressions like (1) in Eq. (2), 
interchange the limit and integration, and Proposition 1 follows.
We have thereby proved a recent Conjecture II put forth by Kreminski
\cite{kreminski}.  

Proposition 2.  The multiplication formula satisfied by the Bernoulli
polynomials \cite{nbs} carries over to $P_n$.  Applying this to the
representation (2) yields the Proposition.

Proposition 3.  One procedure is to take derivatives
of the special case of the Hurwitz zeta function 
$\sum_{r=1}^{q-1} \zeta(s,r/q)=(q^s-1)\zeta(s)$ \cite{hansen} 
and to evaluate as $s \to 1^+$.  Another way is to use
the defining Laurent expansions for the Stieltjes constants 
\cite{kreminski,stieltjes,zhang} in this special case, writing
$$\sum_{r=1}^{q-1}\left [{1 \over {s-1}}+\sum_{k=0}^\infty {{(-1)^k \gamma_k
(r/q)} \over k!} (s-1)^k\right] =\left[qe^{(s-1)\ln q} - 1\right]\zeta (s)$$
$$=\left[q\sum_{j=0}^\infty {{\ln^j q} \over {j!}} (s-1)^j-1\right ]
\left[{1 \over {s-1}}+\sum_{k=0}^\infty {{(-1)^k \gamma_k} \over
k!} (s-1)^k \right].  \eqno(3)$$ 
We expand the right side of this equation and cancel the term $(q-1)/(s-1)$
on both sides.  We then reorder the double sum on the right side and equate
coefficients of like powers of $s-1$ on each side of the resulting equation.
The Proposition follows once again.  Proposition 3 takes the same form as
Corollary 11 of Dilcher \cite{dilcher}, who introduced a generalized
digamma function.  We explain this more in the Discussion section.

Proposition 4.  The key here is to note that 
$$\int_0^1 \zeta(s,a) da = 0, ~~~~~~~~ \sigma < 1,  \eqno(4)$$
and then to integrate the Laurent expansion of the Hurwitz zeta function.
Remark.  Many other related integral-based results for the Stieltjes
constants are possible.

Proposition 5.  For proof, we give the details for $C_{2n-1}(a)$, 
with the case for $C_{2n}(a)$ being very similar.  From Eq. (2) we have
$$C_{2n-1}(a) = (2n-1)!\sum_{k=0}^{2n} {{s(2n,2n-k)} \over {k!}}
\int_1^\infty P_{2n-1}(x-a) {{\ln^k x} \over x^{2n}} dx, ~~n \geq 1. 
\eqno(5)$$
We have \cite{nbs}
$$P_{2n-1}(x-a)=(-1)^n{{2(2n-1)!} \over {(2\pi)^{2n-1}}}\left[\sin 2\pi (x-a)
+ O(2^{1-2n}) \right ], \eqno(6)$$
so that we obtain
$$C_{2n-1}(a)=\kappa_n \sum_{k=0}^{2n} {{s(2n,2n-k)} \over {k!}}\int_1^\infty
\left[\cos 2\pi a\sin 2\pi x-\sin 2\pi a\cos 2\pi x+O(2^{1-2n})\right]
{{\ln^k x} \over x^{2n}} dx,  \eqno(7)$$
where $\kappa_n \equiv (-1)^n2[(2n-1)!]^2/(2\pi)^{2n-1}$.
As a result of the integration in this equation, the leading term in
$C_{2n-1}(a)$ may be written as
$$C_{2n-1}(a) \sim r_1(n)\cos 2\pi a -r_2(n)\sin 2\pi a, \eqno(8)$$
where
$$r_1(n)=\kappa_n \sum_{k=0}^{2n} {{s(2n,2n-k)} \over {k!}}\int_1^\infty
\sin 2\pi x {{\ln^k x} \over x^{2n}} dx,  \eqno(9a)$$
and
$$r_2(n)=\kappa_n \sum_{k=0}^{2n} {{s(2n,2n-k)} \over {k!}}\int_1^\infty
\cos 2\pi x {{\ln^k x} \over x^{2n}} dx.  \eqno(9b)$$
The form (9) can always be written as given in the Proposition with
$\tan 2\pi \phi=-r_1/r_2$ and $m^2 = r_1^2 + r_2^2$.
Remark.  The integrals appearing in Eqs. (9) can be written in several
different ways via integration by parts, but we do not pursue this here.

Proposition 6 follows from Eq. (2), the multiplication formula satisfied
by $P_n$ \cite{nbs}, and the interchange of two finite sums.

Proposition 7.  
According to the property $\partial \zeta(s,a)/\partial a=-s\zeta(s+1,a)$
and the Laurent expansion of $\zeta(s,a)$ we have
$$\sum_{k=0}^\infty {{(-1)^k} \over k!} {{\gamma_k(a)} \over {da}}(s-1)^k
=-1-\sum_{k=0}^\infty {{(-1)^k} \over k!} \gamma_k(a) s^{k+1},
\eqno(10)$$
wherein $d\gamma_0(a)/da=-\psi'(a)$ and $\psi'$ is the trigamma function.
If we perform a binomial expansion on the right side of this equation, 
reorder the sums there, and then equate coefficients of like powers of $s-1$
on both sides, we arrive at the stated set of equations.

Proposition 8.
Valid in the whole complex plane is the form of the Riemann zeta function
$$\zeta(s)={1 \over {1-2^{1-s}}}\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=0}^n (-1)^k {n \choose k} {1 \over {(k+1)^s}}, ~~~~s \neq 1.
\eqno(11)$$
This expression, due to Hasse \cite{hasse}, can be derived by applying
Euler's series transformation to the alternating zeta function \cite{sondow}.

For purposes of expanding Eq. (11) about $s=1$ we have 
$$1-2^{1-s}=-\sum_{j=1}^\infty {{(-\ln 2)^j} \over {j!}} (s-1)^j,
\eqno(12a)$$
$$(k+1)^{-s}=\sum_{\ell=0}^\infty{{(-1)^\ell} \over {\ell!}} \ln^\ell(k+1)
\sum_{q=0}^\ell {\ell \choose q} (s-1)^q, \eqno(12b)$$
and 
$$[1-2^{1-s}]^{-1}={1 \over {\ln 2(s-1)}}+{1 \over 2}+{{\ln 2} \over {12}}
(s-1)- {{\ln^3 2} \over {720}}(s-1)^3+O[(s-1)^5].  \eqno(12c)$$

By using the series $\sum_{k=0}^n {n \choose k} {{(-1)^k} \over {(k+1)}}
={1 \over {n+1}}$ it is easy to see that the $q=0$ term of Eq. (12b)
contributes a $\ln 2$ term in Eq. (11):
$$\sum_{n=0}^\infty {1 \over 2^{n+1}}\sum_{k=0}^n {n \choose k} {{(-1)^k} 
\over {(k+1)}} =\ln 2.  \eqno(13)$$ 
Therefore, comparing with the Laurent expansion of $\zeta(s)$ we obtain 
parts (a) and (b).

Similarly, if we write
$$\ln \zeta(s)=-\ln(1-2^{1-s})+\ln\left[\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=0}^n (-1)^k {n \choose k} {1 \over {(k+1)^s}}\right ]$$
$$=-\ln(s-1)+{1 \over 2}\ln 2(s-1)+{1 \over {24}}(s-1)^2+O[(s-1)^4]$$
$$+\ln\left[\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=0}^n (-1)^k {n \choose k} \sum_{\ell=0}^\infty{{(-1)^\ell} \over 
{\ell!}} \ln^\ell(k+1) \sum_{q=1}^\ell {\ell \choose q} (s-1)^q\right ],
\eqno(14)$$
we find $-\eta_0=\gamma$ and the expression given in part (c).

From Eq. (11) we have
$$\zeta'(s)=-{{\ln 2} \over {(2^{s-1}-1)}}\zeta(s)
-{1 \over {1-2^{1-s}}}\sum_{n=0}^\infty {1 \over 2^{n+1}}
\sum_{k=1}^n (-1)^k {n \choose k} {{\ln(k+1)} \over {(k+1)^s}}, ~~~~s \neq 1.
\eqno(15)$$
Part (d) follows from the well known value
$\zeta'(0)/\zeta(0)=\ln 2\pi$.
Similarly, based upon the well known values of the zeta function at
positive even integers we obtain part (e) for the Bernoulli numbers.

From the functional equation of the zeta function comes the evaluation
for positive integers $n$
$$\zeta'(-2n)=(-1)^n {{(2n)!\zeta(2n+1)} \over {2(2\pi)^{2n}}}.  \eqno(16)$$
From the representation (11) and another globally but more slowly
convergent series for $\zeta(s)$ due to Hasse \cite{hasse} we then have the
identity of part (f).

\medskip
\centerline{\bf Discussion}
\medskip


We may note the special values of the Stirling numbers of the first kind
\cite{nbs} $s(n,0)=\delta_{n0}$ and $s(n,n)=1$.  Therefore, the general
integral representation of Eq. (2) can be simplified to
$$C_n(a) = (-1)^{n-1} n!\left[\int_1^\infty {{P_n(x-a)} \over x^{n+1}}dx
+ \sum_{k=1}^n {{s(n+1,n+1-k)} \over {k!}} \int_1^\infty P_n(x-a) 
{{\ln^k x} \over x^{n+1}} dx \right], ~~n \geq 1. \eqno(17)$$
When $a=1$, the first term on the right side of this equation may be 
evaluated by using integration by parts on the result of the Appendix.

Despite some of the complexity of the representation (11), we have 
several motivations for examining it.  (i)   This form gives rise to
series that play a significant role in renormalization and quantum
field theory \cite{bloch,el}.  
(ii) Equation (11) has connections to random variables in an analytic
number theory setting \cite{biane}.  There is a possibility in particular
to link this structure with that of Brownian processes.
This brings us to a third point of interest.  (iii)  We would like to 
know not only the structure of the Stieltjes constants but also of
the constants $\{\eta_j\}$ that enter the Laurent expansion of the
logarithmic derivative of $\zeta(s)$.  Such knowledge would be very
influential in deciding the nature of a sum denoted $S_2(n)$ 
\cite{coffey,coffey03,coffeygam1} and thereby
the resulting Li/Keiper constants $\lambda_k$ \cite{li,keiper}.

As a byproduct of this work we obtain interesting infinite series for
fundamental constants such as the Euler constant and $\ln \pi$.  The
rapidity of convergence may make some of these suitable for 
applications.  Indeed, even a naive numerical implementation of part (a)
or (b) of Proposition 8 appears to provide these constants to 16 decimal
places after summing over $n$ to $51$ or $52$ terms.

The generalized digamma function used by Dilcher \cite{dilcher} is given by
$$\psi_k(a) \equiv -\gamma_k-{1 \over a}\ln^k a-\sum_{\nu=1}^\infty
\left[{{\ln^k(\nu+a)} \over {\nu+a}}-{{\ln^k a} \over \nu}\right], ~~~~
k \geq 0, \eqno(18)$$
and we may relate it to other functions defined in the literature.
We have $R_m(a)=(-1)^{m+1}(\partial^m/\partial s^m)\zeta(0,a)$ \cite{kanemitsu}
and find that $\psi_k(a)=R'_{k+1}(a)/(k+1)$.  We have
$R_j(a)=(-1)^j j!-\sum_{\ell=0}^\infty \gamma_{j+\ell}(a)/\ell!$ and by using
our Proposition 7 we determine that $\psi_j(a) = -\gamma_j(a)$.  This relation
explains how our Proposition 3 takes exactly the same form as Dilcher's 
Corollary 11 and permits the re-expression of several of his other results as
well.

In his computational work based upon Newton-Cotes integration for the
high accuracy approximation of the Stieltjes constants, Kreminski observed
that for large values of $k$, $\gamma_k(1/2) \approx -\gamma_k$ and that
more generally $C_k(a+1/2) \approx -C_k(a)$.  That is, 
in a sense, the Stieltjes constants for large index are anti-periodic
with period $1/2$.  Our Proposition 1 makes this precise.

Much earlier, Hansen and
Patrick \cite{hansen} showed that a fundamental interval in $a$ for the 
Hurwitz zeta function need only be of length $1/2$.  Perhaps solely on
that basis one would then suspect that some sort of relationship(s) should
exist between $\gamma_k(a)$ and $\gamma_k(a+1/2)$.

Much of our development relies on the underlying theory of periodized Bernoulli
polynomials and corresponding integral representations of the 
Stieltjes constants \cite{zhang}.  Equations such as (17) and still others
that we have obtained help to expose more of the analytic structure of 
the Stieltjes and $\eta_j$ constants. 

\pagebreak
\centerline{\bf Appendix:  Integrals over periodic Bernoulli polynomials}

Herein we evaluate integrals over the periodic Bernoulli polynomial
$P_1$ in terms of polygamma functions $\psi^{(j)}$.  We demonstrate
{\newline \bf Proposition} for positive integers $n$ and $m$
$${{(-1)^n} \over {(n+1)!}} \psi^{(n)}(s)+{1 \over {(n+1)}}\sum_{k=0}^m
{1 \over {(x+k)^{n+1}}}=\int_m^\infty {{P_1(x)} \over {(x+s+1)^{n+2}}}dx
+{1 \over {(n+1)}}{1 \over {(s+m)^{n+1}}}$$
$$-{1 \over {n(n+1)}}{1 \over 
{(s+m+1)^n}}-{1 \over 2}{1 \over {(n+1)}}{1 \over {(s+m+1)^{n+1}}}.  
\eqno(A.1)$$

Our starting point is \cite{edwards}
$$\ln \Gamma(s+1)=(s+1/2)\ln s - s+{1 \over 2}\ln 2\pi-\int_0^\infty
{{P_1(x)} \over {(x+s)}} dx, \eqno(A.2)$$
so that
$$\psi(s+1)=\ln s+ {1 \over {2s}}+\int_0^\infty {{P_1(x)} \over {(x+s)^2}}dx,
\eqno(A.3)$$
where $\psi(s)=\psi(s+1)-1/s$ is the digamma function.  We now take $n$ 
derivatives of relation (A.3), obtaining
$${{(-1)^n} \over {(n+1)!}}\psi^{(n)}(s+1)=-{1 \over {n(n+1)}}{1 \over s^n}
+{1 \over 2}{1 \over {(n+1)}}{1 \over s^{n+1}}+\int_0^\infty {{P_1(x)} \over
{(x+s)^{n+2}}} dx.  \eqno(A.4)$$
Finding that
$$\int_0^\infty {{P_1(x)} \over {(x+s)^{n+2}}} dx =
\int_1^\infty {{P_1(x)} \over {(x+s)^{n+2}}}dx+{1 \over {n(n+1)}}\left[{1 
\over s^n}-{1 \over {(s+1)^n}}\right ]+{1 \over {2(n+1)}}\left[{1 \over s^{n+1}}
-{1 \over {(s+1)^{n+1}}}\right ], \eqno(A.5)$$
we obtain
$${{(-1)^n} \over {(n+1)!}}\psi^{(n)}(s+1)={1 \over {(n+1)}}{1 \over s^{n+1}}
-{1 \over {n(n+1)}}{1 \over {(s+1)^n}}-{1 \over 2}{1 \over {(n+1)}}{1 \over 
{(s+1)^{n+1}}}+\int_1^\infty {{P_1(x)} \over {(x+s)^{n+2}}}dx.  \eqno(A.6)$$

By making a simple change of variable in the integral of Eq. (A.6) and using
the periodicity of $P_1$, we have
$${{(-1)^n} \over {(n+1)!}}\psi^{(n)}(s+m+1)={1 \over {(n+1)}}{1 \over 
{(s+m)^{n+1}}}-{1 \over {n(n+1)}}{1 \over {(s+m+1)^n}}$$
$$-{1 \over 2}{1 \over {(n+1)}}{1 \over {(s+m+1)^{n+1}}}+\int_m^\infty 
{{P_1(x)} \over {(x+s+1)^{n+2}}}dx.  \eqno(A.7)$$
Applying the functional equation of the polygamma function \cite{nbs} we 
obtain the Proposition.

When applying Eq. (A.1), it is useful to keep in mind a relation \cite{grad,sri}
between the polygamma function and the Hurwitz zeta function: $\psi^{(n)}(x)
=(-1)^{n+1}n!\zeta(n+1,x)$.  In particular, when $x=1/2$ or $x=1$ in the
latter relation, values of the Riemann zeta function appear.  In fact,
another starting point for evaluating integrals similar to (A.1) is to
use integral representations for $\zeta(s)$ \cite{nbs,ivic,sri,titch} in terms 
$P_1(x)+1/2$.




\pagebreak

\end{document}